# AdaSwitch: Adaptive Switching Generation for Knowledge Distillation

### v1

@@ Abstract
Small language models (SLMs) are crucial for applications with strict latency and computational constraints, yet achieving high performance remains challenging. Knowledge distillation (KD) can transfer capabilities from large teacher models, but existing methods involve trade-offs: off-policy distillation provides high-quality supervision but introduces a training–inference mismatch, while on-policy approaches maintain consistency but rely on low-quality student outputs. To address these issues, we propose AdaSwitch, a novel approach that dynamically combines on-policy and off-policy generation at the token level. AdaSwitch allows the student to first explore its own predictions and then selectively integrate teacher guidance based on real-time quality assessment. This approach simultaneously preserves consistency and maintains supervision quality. Experiments on three datasets with two teacher–student LLM pairs demonstrate that AdaSwitch consistently improves accuracy, offering a practical and effective method for distilling SLMs with acceptable additional overhead.
AdaSwitch: Adaptive Switching Generation for Knowledge Distillation
@@ 1 Introduction
With the rapid development of large language models (LLMs), scaling laws suggest that increasing the number of parameters generally leads to better performance. However, small language models (SLMs) remain crucial in scenarios where latency and computational cost are highly sensitive. For instance, in search engines and recommender systems, it is common to process a large volume of queries simultaneously Li et al. (2025); Deng et al. (2025). Although LLMs may achieve higher accuracy, their high latency and computational demands can be prohibitive in such real-world applications. Moreover, SLMs also play an important role in accelerating inference methods such as speculative decoding Leviathan et al. (2023); Xia et al. (2024); Yin et al. (2024). Therefore, developing an SLM that achieves strong performance on a specific task is of vital importance.
@@
Figure 1: An example illustrating the sequence generation process for off-policy, on-policy, and SKD methods. Refer to caption denotes the teacher, and Refer to caption represents the student.
Obtaining high-quality SLMs is often achieved by transferring knowledge from large models. Among various knowledge transfer techniques, knowledge distillation (KD) has emerged as a particularly effective approach Xu et al. (2024); Wang et al. (2025); Sanh et al. (2019). Unlike other transfer methods that may require extensive retraining on annotated datasets or rely solely on task-specific supervision, KD leverages the outputs of a teacher model to guide the student model Hinton et al. (2015). By utilizing the teacher’s predictions, including intermediate representations or logits, KD provides richer soft labels that capture nuanced information beyond hard ground-truth labels Yu et al. (2024); Aguilar et al. (2020). This enables the student to better approximate the teacher’s behavior, often leading to improved generalization, robustness, and efficiency, especially in scenarios with limited annotated data.
In KD, the selection of target sequences for sampling and supervision plays a pivotal role, as it directly shapes the quality and distribution of the learning signal. Based on this, KD methods are often categorized as off-policy or on-policy depending on the source of training data Xu et al. (2024). Off-policy approaches update the student model using either ground-truth sequences as hard labels or the teacher’s logits over sequences generated by the teacher Kim and Rush (2016). While this strategy ensures relatively high-quality training sequences, it can lead the student to suffer from a training–inference mismatch, as inference relies solely on the student’s own outputs. To address this issue, on-policy methods use the teacher’s logits on sequences generated by the student, thereby aligning the training process more closely with inference Agarwal et al. . However, due to the student’s limited capabilities, the quality of these student-generated sequences can be poor, especially in the early stages of distillation, restricting the effectiveness of distillation Xu et al. . Consequently, effectively combining the strengths of on-policy and off-policy approaches is key to unlocking the full potential of KD.
Most existing work focuses on data-level combinations, where sequences generated by the student and the teacher are simply mixed Lin et al. (2020); Agarwal et al. . However, this coarse-grained strategy cannot guarantee both high quality and consistency for each individual sequence. At the token-level, the only notable attempt is SKD, which introduces speculative decoding into sequence generation: the teacher supervises the student’s output at each token step Xu et al. . While this can align generation more closely with the teacher, the frequent switching between models produces “twisted” sequences that overfit to the teacher, yielding only marginal performance gains.
To address these limitations, we propose AdaSwitch, a two-stage sequence generation framework that dynamically alternates between on-policy and off-policy learning at the token level. In the exploration stage, the student generates an initial sequence autonomously, maintaining training–inference consistency while exploring with its latent capabilities. In the guidance stage, the teacher takes over and generates the remaining sequence to ensure high-quality outputs. Furthermore, the switching threshold is dynamically adjusted during generation, allowing the switch to adapt based on the difficulty of each sample. This one-time switch mechanism prevents excessive interventions, reduces sequence distortion and teacher overfitting, and maintains a balance between consistency and quality.
@@
2 Methodology
2.1 Problem Setup
$D(M_{t}\|M_{s})(y|x)=\frac{1}{L_{y}}\sum_{i=1}^{L_{y}}d_{i}(x,y),$ (2)
2.2 Target Sequence Selection
$y\sim M_{s}(\cdot\mid x;\theta_{s}).$ (3)
@@ 2.3 AdaSwitch
Llama-3.1-3b to Llama-3.1-1b Qwen-2.5-3b to Qwen-2.5-0.5b
SUMM GSM GSM_Plus SUMM GSM GSM_Plus
Teacher 37.45 59.59 38.58 37.41 65.35 45.42
Student 13.48 0.00 0.00 12.07 0.00 0.00
SFT 32.34 29.04 12.00 30.28 31.01 15.54
SeqKD 32.09 28.20 11.78 30.13 30.42 14.42
Off-policy KD 34.57 30.55 13.83 32.44 33.51 17.92
On-policy KD 34.34 32.68 14.33 32.31 34.12 20.17
ImitKD 35.07 30.55 15.63 32.50 34.57 21.29
SKD 34.88 32.37 16.88 32.27 32.98 20.38
AdaSwitch 35.56 35.03 17.33 32.60 36.32 21.63
@@ 3.1 Student and Teacher Models
We conducted experiments using two state-of-the-art LLM families: Qwen 2.5 Team (2024) and Llama 3.1 Grattafiori et al. (2024). For the teacher models, we employed the SFTed Qwen-2.5-3B and Llama-3.1-3B, while the student models consisted of Qwen-2.5-0.5B and Llama-3.1-1B.
@@ 3.3 Hyperparameters
For all SFT processes, we used a learning rate of 1e-5, a warmup ratio of 0.1, and a dropout rate of 0.1. Each SFT checkpoint was trained for three epochs, with the checkpoint exhibiting the lowest validation loss being selected. For the KD process, we set the maximum input and output lengths to {1024, 256} for dialogue summarization and {128, 512} for arithmetic reasoning. The student models were configured with a temperature of 0.5 and a top-p of 0.5, while the teacher models had a temperature of 0.2 and a top-p of 0.5, as per previous research. All tasks and models were trained for three epochs with a learning rate of 1e-5. To ensure a fair comparison, the $K_{skd}$ for SKD was set to 0.5, as reported in its original paper. We set the sliding window length $L$ to 10 and the threshold $K$ to 3 without specific fine-tuning. The impact of these hyperparameters is discussed later.
@@ 3.4 Baselines
Supervised FT directly minimize the negative log-likelihood of the student’s predictions on the target sequence, which is the ground-truth.
@@ 3.5 Overall Performance
Mixed methods such as ImitKD and SKD attempt to integrate both paradigms and yield localized improvements. For instance, ImitKD performs well on SUMM, and SKD shows gains on GSM-Plus. However, these improvements are marginal and inconsistent across tasks and model pairs, reflecting the limited performance improvement of their designs.
@@
3.6 Analysis on Switch Rate and KLD
@@
First, we observe a significant decrease in KLD during distillation across all tasks. This trend indicates that the student model gradually converges toward the teacher, consistent with intuitive expectations. Moreover, the KLD in the SUMM task is consistently lower than that in GSM, aligning with the performance difference between the two tasks and supporting the view that SUMM is relatively easier than GSM.
@@
Further, we observe that the switch rate increases steadily during distillation, most notably in SUMM, while remaining near 100% in GSM. This trend reflects a gradual decline in exploration and a rise in guidance, indicating that on-policy distillation contributes less over time. Although this may seem counterintuitive, since weaker students would normally require more guidance, the pattern is explained by KLD: early in training, high divergence suppresses switching, but as the student aligns with the teacher, switching becomes more frequent. In other words, AdaSwitch begins in an exploration-dominated on-policy mode and gradually shifts to off-policy guidance as the student improves. This adaptive behavior allows AdaSwitch to dynamically balance exploration and guidance according to the student’s capability, ensuring optimal distillation across varying tasks. Due to space limitations, we provide the complete experimental results and the analysis of relative switch positions during distillation in Appendix B.
@@ 3.7 Early-Stage Distillation Analysis
Further experimental results of the performance on the test set reveals that AdaSwitch demonstrates the most rapid improvement rate, particularly on the GSM dataset. Notably, at step 30, the accuracy of AdaSwitch surpasses 10%, whereas the off-policy and on-policy methods achieve only 7.96% and 0%, respectively. These results suggest that AdaSwitch, through effective token-level mixing, simultaneously ensures the quality of generated sequences and reduces the discrepancy between training and inference sequences during the early stages of distillation. The full experimental results are presented in Appendix B.
@@ 3.9 Time Consumption Analysis
In AdaSwitch, the time cost primarily depends on the sequence length and the per-token generation speed. However, in practice, it is difficult to estimate the sequence length for each sample across different KD methods, since generation is performed by different models. Moreover, sequence generation accounts for only a small fraction of KD time. Therefore, we conducted a general and practical experiment by comparing the total runtime of different KD methods over the distillation.
The results in Table 2 indicate that off-policy consistently achieves the lowest runtime cost, since it doesn’t conduct sequence generation. On the other hand, both SKD and AdaSwitch are slower than on-policy because they require computing teacher logits during generation and switching between the student and teacher.
@@
Figure 6: Parameter analysis on $K$ and $L$.
@@ 4.1 Small Language Model Optimization
4.2 Knowledge Distillation for LLMs
In the era of LLM, KD has been adapted to autoregressive models and categorized into off-policy and on-policy methods, inspired by reinforcement learning Xu et al. (2024). Off-policy KD uses either ground-truth sequences or teacher-generated logits for supervision Kim and Rush (2016); Sanh et al. (2019). While this ensures high-quality training data, it often causes a mismatch between the student’s training distribution and inference-time behavior, leading to inconsistencies Agarwal et al. . Conversely, on-policy KD trains the student using its own generated sequences, improving consistency and performance on some tasks. However, the student’s initially poor outputs limit sequence quality and thus KD effectiveness Xu et al. . Prior attempts to combine off- and on-policy KD typically mix data at a coarse level Lin et al. (2020); Agarwal et al. . The most related approach, SKD, employs speculative decoding guided by teacher supervision to enhance sequence quality Xu et al. . However, its frequent switching between student and teacher models during generation can cause overfitting to the teacher and restrict performance improvements. In contrast, AdaSwitch mitigates this issue by introducing a more stable and efficient token-level mixing strategy.
@@ 5 Conclusion
In this paper, we introduce AdaSwitch, a token-level mixed distillation framework designed to reduce the gap between training and inference while preserving high-quality student outputs. AdaSwitch adopts a two-stage generation process that dynamically switches from student exploration to teacher guidance. Extensive experiments on three tasks, two LLM families, and three distance metrics verify its effectiveness, delivering significant performance gains with acceptable time cost.
@@ 6 Limitations
Our primary limitation lies in the constrained scope of our experimental evaluation. We validate AdaSwitch on three tasks and two LLM families, which may not fully capture the breadth of challenges faced in diverse real-world applications. Future work will extend the evaluation to a wider range of tasks and domains to further verify the generality and robustness of the proposed method. Additionally, while AdaSwitch achieves significant performance gains, it introduces an additional computational overhead during the KD process. As shown in our time complexity analysis, AdaSwitch incurs approximately 20% higher time consumption compared to the standard on-policy KD method. However, we consider this trade-off acceptable because the KD process is typically performed offline, and crucially, our method introduces no additional latency or computational cost during the inference phase for the final student model.
@@ Appendix B More Experimental Results
Figure 9: Analysis of the relative switching position of distillation process

### v3

# AdaSwitch: Balancing Exploration and Guidance in Knowledge Distillation via Adaptive Switching

@@ Abstract
Small language models (SLMs) are crucial for applications with strict latency and computational constraints, yet achieving high performance remains challenging. Knowledge distillation (KD) can transfer capabilities from large teacher models, but existing methods face a dilemma: off-policy distillation provides high-quality supervision but suffers from exposure bias (training–inference mismatch), while on-policy approaches ensure consistency but are limited by the low quality of student-generated outputs. To address these issues, we propose AdaSwitch, a novel approach that dynamically combines on-policy and off-policy generation via an adaptive switching mechanism. AdaSwitch allows the student to explore its predictions within its capability and selectively integrates teacher guidance only when divergence exceeds a context-aware threshold. This paradigm preserves generation consistency while ensuring high-quality supervision. Experiments on three datasets demonstrate that AdaSwitch consistently improves accuracy and reasoning capability with moderate overhead.
AdaSwitch: Balancing Exploration and Guidance in Knowledge Distillation via Adaptive Switching
@@ 1 Introduction
With the rapid development of large language models (LLMs), scaling laws suggest that increasing parameters generally leads to better performance. However, in latency-sensitive scenarios such as real-time search and recommendation systems, deploying massive LLMs is often computationally prohibitive Li et al. (2025); Deng et al. (2025). Therefore, developing Small Language Models (SLMs) that retain strong performance while minimizing inference costs is of vital importance.
@@
Figure 1: An example illustrating the sequence generation process for off-policy, on-policy, SKD, and SWITCH methods. Refer to caption denotes the teacher, and Refer to caption represents the student.
Obtaining high-quality SLMs is often achieved by transferring knowledge from large teacher models via Knowledge Distillation (KD) Hinton et al. (2015); Xu et al. (2024); Wang et al. (2025); Sanh et al. (2019). Current KD paradigms generally fall into two categories depending on the source of training data Xu et al. (2024): Off-policy methods (e.g., MiniLLM  Gu et al. (2024), SeqKD  Kim and Rush (2016)) utilize teacher-generated sequences or ground-truth data, providing rich supervision signals but suffering from training-inference mismatch, as the student never learns to recover from its own errors. Conversely, on-policy methods (e.g., GKD  Agarwal et al. ) train on student-generated sequences, aligning training with inference but often collapsing due to the low quality of the student’s initial outputs, especially in the early stages of distillation. Effectively combining the strengths of both paradigms—high-quality supervision and distribution consistency—remains a key challenge.
Recent attempts, such as SKD Xu et al. and SWITCH Koo et al. (2025), propose mixing student and teacher tokens within a single sequence. However, these methods typically employ frequent, bidirectional switching strategies, where control is returned to the student immediately after a teacher intervention. We argue that such fragmentation disrupts the semantic coherence of the generation process. When a model is forced to constantly toggle between the student’s and the teacher’s latent states, it may generate disjointed sequences (or "chimeras") that satisfy neither distribution perfectly, as validated in Section 3.9. Furthermore, existing static or rigid thresholding mechanisms fail to adapt to the varying difficulty of different samples, limiting their effectiveness on complex reasoning tasks.
To address these limitations, we propose AdaSwitch, a context-aware sequence generation framework that dynamically alternates between on-policy exploration and off-policy guidance at the token level. AdaSwitch adopts an adaptive one-time switching mechanism: the student initially generates tokens autonomously (Exploration Stage) to maintain training-inference consistency. Once the student’s divergence from the teacher exceeds a dynamic threshold derived from recent generation history, the teacher takes over to generate the remaining sequence (Guidance Stage), ensuring high-quality outputs. This design prevents logic fragmentation caused by frequent switching and ensures that the student receives a high-quality correction exactly when its capability limit is reached.
@@
2 Methodology
2.1 Problem Setup
$D(M_{t}\|M_{s})(y|x)=\frac{1}{L_{y}}\sum_{i=1}^{L_{y}}d_{i}(x,y),$ (2)
2.2 Target Sequence Selection
$y\sim M_{s}(\cdot\mid x;\theta_{s}).$ (3)
Llama-3.1-3b to Llama-3.1-1b Qwen-2.5-3b to Qwen-2.5-0.5b Gemma-7b to Gemma-2b
SUMM GSM GSM_Plus SUMM GSM GSM_Plus SUMM GSM GSM_Plus
Teacher 37.5 59.6 38.6 37.4 65.4 45.4 57.0 36.9 33.4
Student 13.5 0.0 0.0 12.1 0.0 0.0 11.6 0.0 0.0
SFT 32.3 29.0 12.0 30.3 31.0 15.5 31.8 18.9 11.7
SeqKD 32.1 28.2 11.8 30.1 30.4 14.4 31.0 21.8 10.5
Off-policy 34.6 30.6 13.8 32.4 33.5 17.9 35.1 22.6 10.7
On-policy 34.3 32.7 14.3 32.3 34.1 20.2 34.1 26.2 13.8
ImitKD 35.1 30.6 15.6 32.5 34.6 21.3 34.9 25.3 13.5
SWITCH 33.7 33.2 16.0 31.8 35.2 20.8 35.1 24.8 13.8
SKD 34.9 32.4 16.9 32.3 33.0 20.4 35.0 26.0 13.9
AdaSwitch 35.6 35.0 17.3 32.6 36.3 21.6 35.6 26.8 14.2
@@ 3.1 Student and Teacher Models
We conducted experiments on three state-of-the-art LLM families, following prior work  Agarwal et al. ; Koo et al. (2025); Xu et al. in selecting comparable teacher–student model scales: Qwen 2.5 Team (2024), Llama 3.1 Grattafiori et al. (2024), and Gemma Team et al. (2024). For the teacher models, we employed the SFTed Qwen-2.5-3B, Llama-3.1-3B and Gemma-7B, while the student models consisted of Qwen-2.5-0.5B, Llama-3.1-1B and Gemma-2B.
@@ 3.3 Hyperparameters
For all SFT experiments, we used a learning rate of 1e-5, a warmup ratio of 0.1, and a dropout rate of 0.1, training each model for three epochs and selecting the checkpoint with the lowest validation loss. For KD, input/output lengths were set to 1024, 256 for dialogue summarization and 128, 512 for arithmetic reasoning. Student models used a temperature of 0.5 and top-p of 0.5, while teachers used a temperature of 0.2 and top-p of 0.5, following prior work Xu et al. . All models were trained for three epochs with a learning rate of 1e-5. For fair comparison, we set $K_{skd}=0.5$ for SKD, and fixed the sliding window length $L=10$ and threshold $K=3$. The effects of $L$ and $K$ are analyzed later.
@@ 3.4 Baselines
We evaluate the effectiveness of AdaSwitch by comparing it against six baseline methods, spanning pure on-policy and off-policy approaches, and existing sequence-level and token-level mixing methods: Supervised FT, SeqKD Kim and Rush (2016), SupervisedKD Sanh et al. (2019), GKD Agarwal et al. , ImitKD Lin et al. (2020), SWITCH Koo et al. (2025), and SKD Xu et al. . Due to space constraints, detailed descriptions of these baselines are provided in Appendix A.1.
@@ 3.5 Overall Performance
Mixed methods such as ImitKD, SWITCH and SKD attempt to integrate both paradigms and yield localized improvements. For instance, ImitKD performs well on SUMM, and SKD shows gains on GSM-Plus. However, these improvements are marginal and inconsistent across tasks and model pairs, reflecting the limited performance improvement of their designs.
@@
Figure 4: Analysis of the relative switching position of the distillation process
3.6 Analysis of Switching Dynamics during Distillation
@@
First, we observe a significant decrease in KLD during distillation across all tasks. This trend indicates that the student model gradually converges toward the teacher, consistent with intuitive expectations. Moreover, the KLD in the SUMM task is consistently lower than that in GSM, aligning with the performance difference between the two tasks and indicating that the SUMM task presents a lower reasoning complexity compared to GSM.
@@
Further, we observe that the switch rate increases steadily during distillation, most notably in SUMM, while remaining near 100% in GSM. This trend reflects a gradual decline in exploration and a rise in guidance, indicating that on-policy distillation contributes less over time. Although this may seem counterintuitive, since weaker students would normally require more guidance, the pattern is explained by KLD: early in training, high divergence suppresses switching, but as the student aligns with the teacher, switching becomes more frequent. In other words, AdaSwitch begins in an exploration-dominated on-policy mode and gradually shifts to off-policy guidance as the student improves. This adaptive behavior allows AdaSwitch to dynamically balance exploration and guidance according to the student’s capability, ensuring optimal distillation across varying tasks.
@@ 3.7 Early-Stage Distillation Analysis
Further experimental results of the performance on the test set reveal that AdaSwitch demonstrates the most rapid improvement rate, particularly on the GSM dataset. Notably, at step 30, the accuracy of AdaSwitch surpasses 10%, whereas the off-policy and on-policy methods achieve only 7.96% and 0%, respectively. These results suggest that AdaSwitch, through effective token-level mixing, simultaneously ensures the quality of generated sequences and reduces the discrepancy between training and inference sequences during the early stages of distillation. The full experimental results are presented in Appendix B.
@@ 3.8 Parameter Analysis
Figure 7: Parameter analysis on $K$ and $L$.
3.9 Impact of Maximum Switching Frequency
To validate our single-switch design, we conducted an ablation study varying the maximum switching budget from 1 to 3 on SUMM and GSM tasks using the Llama model pair. Averaged over five random seeds, the results in Figure 8 show that increasing the switch limit consistently degrades performance: Rouge-L on SUMM drops from 35.6 to 35.2, while GSM accuracy declines from 35.0 to 34.6. Beyond performance loss, we observe marked training instability, with error bars expanding significantly as the budget grows, particularly on GSM. This suggests that frequent student–teacher alternations introduce optimization noise and fragment the generation process into semantically inconsistent “chimera” sequences. These findings empirically justify AdaSwitch’s one-time switching mechanism, which triggers at most once per sequence precisely at the student’s capability limit, thereby preserving generation coherence while ensuring timely guidance.
@@ 3.10 Time Consumption Analysis
Figure 8: Performance under maximum number of switches allowed
@@ 4.1 Small Language Model Optimization
4.2 Knowledge Distillation for LLMs
In the era of LLM, KD has been adapted to autoregressive models and categorized into off-policy and on-policy methods, inspired by reinforcement learning Xu et al. (2024). Off-policy KD uses either ground-truth sequences or teacher-generated logits for supervision Kim and Rush (2016); Sanh et al. (2019). While this ensures high-quality training data, it often causes a mismatch between the student’s training distribution and inference-time behavior, leading to inconsistencies Agarwal et al. . Conversely, on-policy KD trains the student using its own generated sequences, improving consistency and performance on some tasks. However, the student’s initially poor outputs limit sequence quality and thus KD effectiveness Xu et al. . Prior attempts to combine off- and on-policy KD typically mix data at a coarse level Lin et al. (2020); Agarwal et al. . Among the two most related approaches, SKD employs speculative decoding guided by teacher supervision to enhance sequence quality Xu et al. . SWITCH Koo et al. (2025) alternates between the student and the teacher during sequence generation based on a threshold with a fixed decay rate. However, both methods rely on frequent switching between the student and teacher models during generation, which can lead to overfitting to the teacher and limit performance improvements. In contrast, AdaSwitch mitigates this issue by introducing a more stable and efficient token-level mixing strategy.
@@ 5 Conclusion
In this paper, we introduce AdaSwitch, a token-level mixed distillation framework designed to reduce the gap between training and inference while preserving high-quality student outputs. AdaSwitch adopts a two-stage generation process that dynamically switches from student exploration to teacher guidance. Extensive experiments on three tasks, three LLM families, and three distance metrics verify its effectiveness, delivering significant performance gains with acceptable time cost.
@@ 6 Limitations
Our primary limitation lies in the constrained scope of our experimental evaluation. We validate AdaSwitch on three tasks and three LLM families, which may not fully capture the breadth of challenges faced in diverse real-world applications. Future work will extend the evaluation to a wider range of tasks and domains to further verify the generality and robustness of the proposed method. Additionally, while AdaSwitch achieves significant performance gains, it introduces an additional computational overhead during the KD process. As shown in our time complexity analysis, AdaSwitch incurs approximately 30% higher time consumption compared to the standard on-policy KD method. However, we consider this trade-off acceptable because the KD process is typically performed offline, and crucially, our method introduces no additional latency or computational cost during the inference phase for the final student model.
@@ Appendix A Datasets
A.1 Baselines
Supervised FT directly minimizes the negative log-likelihood of the student’s predictions on the target sequence, which is the ground-truth.
SWITCH Koo et al. (2025) employs token-level switching between the student and teacher based on a fixed threshold decay during sequence generation.